OpenStudy (anonymous):

9 3/4 percent converted to a fraction equals

OpenStudy (anonymous):

9 3/4% = (9 3/4)/100 = 39/400

OpenStudy (anonymous):

39 39 1 39 --- = ----- x ----- = ------ 4 4 100 400 _____ 100

OpenStudy (anonymous):

9 3/4% = 9.75% = 0.0975